Using gauss theorem calculate the flux of the vector field A = x3i^ + x2zj^ + yzk^ through the surface of a cube of side 2 units

xUsing gauss theorem calculate the flux of the vector field A =


1
Expert's answer
2021-03-01T11:51:05-0500
"div{\vec{A}}=3x^2+y\\\Phi=\int_0^2\int_0^2\int_0^2(3x^2+y)dxdydz \\
div{\vec{A}}=3x^2+y\\\Phi=2^2\int_0^2(3x^2)dx+2^2\int_0^2(y)dy\\\Phi=2^2(2^3)+2^2(0.5\cdot2^2)\\\Phi=40"